OpenStudy (anonymous):

PROVE:

OpenStudy (anonymous):

\[\frac{ 1+\sin \theta }{ 1-\sin \theta }-\frac{ 1-\sin \theta }{ 1+\sin \theta }=4\tan \theta*\sec \theta\]

hartnn (hartnn):

what have u tried ?

OpenStudy (anonymous):

I am stuck at some place

hartnn (hartnn):

which step ?

OpenStudy (anonymous):

-2cos^2theta/cos^2theta

hartnn (hartnn):

denominator, obviously is cos^2 theta lets see numerator from beginning, (1+s)^2-(1-s)^2 {s= sin theta} can you simplify this ?

OpenStudy (anonymous):

1+s/1-s - 1-s/1+s

hartnn (hartnn):

not 1+s^2 its (1+s)(1+s) = (1+s)^2

OpenStudy (anonymous):

(1+s)^2-(1-s)^2/(1-s)(1+s)

hartnn (hartnn):

yes, go ahead

OpenStudy (dan815):